# Multi-objective Representation for Numbers in Clinical Narratives Using CamemBERT-bio

### v2

@@ Abstract
The processing of numerical values is a rapidly developing area in the field of Large Language Models (LLMs). Despite numerous advancements achieved by previous research, significant challenges persist, particularly within the healthcare domain. This paper investigates the limitations of LLMs in understanding numerical values. Objective: The aim of this research is to categorize numerical values extracted from medical documents into eight specific physiological categories using CamemBERT-bio. Methods: Building on prior findings that reveal potential limitations of transformer-based models, we examine two strategies: fine-tuning CamemBERT-bio on a small medical dataset with the integration of Label Embedding for Self-Attention (LESA), and combining LESA with additional enhancement techniques such as Xval. Given that CamemBERT-bio is already pre-trained on a large medical dataset, the first approach aims to update its encoder with newly added label embeddings technique, while the second approach seeks to develop multiple representations of numbers (contextual and magnitude-based) to achieve more robust number embeddings. Results: As anticipated, fine-tuning the standard CamemBERT-bio on our small medical dataset did not improve F1 scores. However, significant improvements were observed with CamemBERT-bio + LESA, resulting in an over 15% increase. Similar enhancements were noted when combining LESA with Xval, outperforming conventional and previous methods. Conclusions and Novelty: This study introduces two innovative techniques for handling numerical data, which are also applicable to other modalities. We illustrate how these techniques can improve the performance of Transformer-based models, achieving more reliable classification results even with small datasets.
Index Terms:
Clinical Natural Language Processing, Numerical values classification, Language models training.
@@ I Introduction
This paper contributes to a broader project aimed at leverage Pretrained Large Language Models (PLMs) for diagnosing cardiac failure. According to the National Heart, Lung, and Blood Institute (2023), cardiac failure occurs when the heart is unable to pump sufficient blood to meet the body’s needs, either due to inadequate filling or diminished pumping capacity. Cardiac failure presents through a range of physiological indicators, including blood pressure, ventricular gradients, and ejection fractions. Our research focuses on classifying numerical values extracted from medical notes into one of eight specific physiological categories using CamemBERT-bio [1].
I-A Goal statement
@@ I-B Contribution
  • We assess and juxtapose our tailored model against conventional counterparts. Findings demonstrate that our model surpasses standard approaches and adeptly manages datasets characterized by limited size and imbalance.
II Related works
@@
Processing numerical values with LLMs is a highly active research area. While [4] claimed that most word embedding models could capture numeracy, several limitations persist. Various approaches have been developed to handle numerical values. [5] tackled the issue of number classification by employing a rules-based method utilizing regular expressions. Nevertheless, the success of this approach is largely contingent on the clarity of the text. It depends significantly on consistent abbreviations and the minimal presence of irrelevant textual noise, which can disrupt the regular expression generator’s accuracy. [6] proposed a question-answer pre-training task aimed at improving number understanding. Other studies have explored the potential of different number tokenization methods, such as digit-by-digit, scientific notation, and customized notations [7]. Some previous studies have tackled the challenge with a number-agnostic approach, replacing numbers with placeholder keywords and scaling their embeddings to enforce context-centric encoding [3, 8, 2]. In their work, [2] introduced Label Embedding for Self-Attention (LESA), which incorporates class descriptions directly into the initial feature extraction phase to help the feature extractor focus on relevant features for classification. LESA is similar to EmbedNum [9], which classifies numbers by comparing their embeddings to those of potential categories. However, LESA differs by using similarity exclusively to construct the number’s embedding rather than for the classification process itself.
@@
Another significant challenge is handling small and imbalanced datasets. While transformer-based architectures generally perform well with large datasets, research by [10, 11, 12] indicates that they may not outperform traditional methods, such as LSTM, when dealing with smaller datasets of clinical notes. A common technique to enhance LLM performance on small datasets is Knowledge Distillation (KD) [13]. Knowledge Distillation (KD) is based on the principle that a model with fewer parameters is less prone to overfitting. This technique entails training a smaller neural network to replicate the performance of a larger one, using a teacher-student framework. This methodology is the key idea of like DistilBERT [14] and its French counterpart, DistillCamemBERT [15]. Another approach to handling limited data is to create multiple representations for words, leveraging diverse methods to achieve more robust and generalizable modeling. Typically, this involves constructing an objective loss as the sum of various losses, each specialized in different aspects of the studied modality. However, this approach raises the critical issue of appropriately weighting each loss to avoid bias. In [16], the authors propose an automatic method for weighting multiple losses by optimizing the log-likelihood of the model predictions. Although their approach was applied to images, we found their results relevant and potentially beneficial for our case.
III The task and the dataset
@@
IV A few limitations of end-to-end learning for LLM in medical datasets
End-to-end learning involves training a model to perform a task from raw input to final output, without intermediate steps or feature engineering. While this is the standard approach for training most LLMs, our work faces the significant limitation of working with small and imbalanced datasets, unlike many state-of-the-art approaches mentioned earlier. Additionally, there are several other major limitations of end-to-end training for LLMs in the medical domain.
IV-A The limitations of Tokenization
@@ IV-B The structure issue
As noted in [2], the last excerpt underscores the lack of contextual information in sentence structures. This excerpt reads a sequence of physiological indicators. The initial reference to FR most likely indicates the respiratory rate (fréquence respiratoire) since a value of 21 is usually too low for the shortening fraction (fraction de raccourcissement). Conversely, the second occurrence of FR refers to a value above 50, suggesting it pertains to the shortening fraction. In this case, it was the values themselves, rather than the context, that provided the necessary distinctions, highlighting a flaw in the sentence structure. This issue echoes the challenges of tokenization, emphasizing the need for a thorough understanding of the numerical values. Additionally, It’s worth mentioning that CamemBERT-bio underwent no pre-training using unclear or poorly structured medical documents.
V Methodology
The methodology adopted in this study involves developing a multi-objective representation of numbers to achieve a more generalizable and robust model [20]. Our objectives are Masked Language Modelling (MLM) loss and Number Prediction Loss. These two objectives result in the derivation of two models:
@@ V Methodology
V-A Masked Language Modelling (MLM) objective
@@
For every individual class, some descriptive keywords are selected, and their initial embeddings are computed and then averaged. This calculation produces an embedding matrix, represented as $\boldsymbol{X}^{l}\in\mathbb{R}^{n\times D}$, where $n$ corresponds to the number of classes. This matrix includes unique label embeddings for each class.
@@
In this work, we train this architecture using the Masked Language Modelling (MLM) task as a prefinetuning step. The MLM task involves randomly masking tokens and asking the model to predict the masked tokens, thereby training the model to build contextual representations of tokens. To address the limited size of our dataset, we employ the LESA mechanism to accelerate the learning process. Essentially, we inject additional information into the model via label-related keywords, enabling the model to more quickly grasp the relationships between tokens.
V-B Number Prediction objective
This method consists in four steps as detailed in [3]. Let’s consider an input text $t$ containing both numbers and words. First, all the numbers are replaced by a placeholder keyword NUM, forming a new input text $t_{masked}$. The initial numbers are saved for later. Then, we tokenize and embed $t_{masked}$ resulting in the matrix $h_{masked}$. Subsequently, all the vectors in $h_{masked}$ corresponding to the numbers previously replaced are multiplied by their corresponding value, resulting in a new matrix $h$. Finally, this embeddings matrix $h$ is fed to the encoder of our model.
By replacing all numbers by a keyword placeholder, we solve the tokenization limitation mentionned earlier. This loss objective aims at training the model to learn the magnitude of the numbers it encodes. The magnitude information is very important as it is a good discriminator between classification classes. For instance the numbers belonging to heart rate class are usually bigger than those belonging to APGAR class.
V-C Multi Objective Loss
Working with a multi objectives losses require to merge them into one single loss
@@ V-C Multi Objective Loss
where $w_{i}$ represents the weight of the loss $L_{i}$. In [16], the authors propose an automatic way to weight multiple losses by optimizing the model predictions log-likelihood. In plain terms, for the number prediction task, the authors assume that the likelihood is given as a Gaussian
$p\left(y|f_{1}(x,w)\right)=\mathcal{N}\left(f_{1}(x,w)|\sigma_{1}^{2}\right)$ (3)
where $x$ is the input sequence, $w$ the model parameters and $f_{1}(x,w)$ the model’s number prediction. $\mathcal{N}$ refers to the gaussian probability distribution function and $\sigma_{1}$ its standard deviation. This leads to the log likelihood
$\log p(y|f_{1}(x,w))\propto-\frac{1}{2\sigma_{1}^{2}}\left(y-f_{1}(x,w)\right)^{2}-\log\sigma_{1}$ (4)
Then one can recognize the square error $\left(y-f_{1}(x,w)\right)^{2}$ with its weight $\frac{1}{2\sigma_{1}^{2}}$ and a regularization term $\log\sigma_{1}$.
This appraoch also works for the MLM task. We write the likelihood as
$p(z|f_{2}(x,w),\sigma_{2})=\textrm{Softmax}\left(\frac{1}{\sigma_{2}^{2}}f_{2}(x,w)\right)_{z}$ (5)
@@
where $f_{2}(x,w)$ is the predicted token distribution vector and $\sigma_{2}$ is artificially added as a temperature in a Boltzmann distribution. Then, after a few approximations, we get to write the log likelihood as
@@
$\log p(z|f_{2}(x,w))\propto\frac{1}{\sigma_{2}^{2}}\log\textrm{Softmax}\left(f_{2}(x,w)\right)_{z}-\log\sigma_{2}$ (6)
@@
Here, one may recognize the cross entropy loss with its weight $\frac{1}{\sigma_{2}^{2}}$ and a regularization term $\log\sigma_{2}$. When we sum up everything, for a given number $y$, we consider the following loss:
@@
$\begin{split}\mathcal{L}(y)&=\frac{1}{2\sigma_{1}^{2}}\left(y-f_{1}(x,w)\right)^{2}-\frac{1}{\sigma_{2}^{2}}\log\textrm{Softmax}\left(f_{2}(x,w)\right)_{y}\\ &+\log\sigma_{1}+\log\sigma_{2}\\ &=\frac{1}{2\sigma_{1}^{2}}\textrm{SEL}+\frac{1}{\sigma_{2}^{2}}\textrm{MLM}_{loss}+\log\sigma_{1}+\log\sigma_{2}\end{split}$ (7)
where SEL stands for the Square Error Loss, that is $\left(y-f_{1}(x,w)\right)^{2}$, and $\textrm{MLM}_{loss}$ is the language modeling loss, that is $\log\textrm{Softmax}\left(f_{2}(x,w)\right)_{y}$.
Now we optimize this lost with respect to $w,\sigma_{1},\sigma_{2}$. In our case, when writing the first-order condition with respect to $\sigma_{1}$, we get
$\sigma_{1}=\left(y-f_{1}(x,w)\right)$ (8)
Since the exact numbers within a medical note are often not predictable from the context, we anticipate that the regression loss $(y-f_{1}(x,w))^{2}$ will have a significant lower bound that we will note $l$. Additionally, this lower bound increases with the widening range of possible values, which is particularly high in our case due to the broad spectrum of numbers. Consequently, at equilibrium, Equation  8 implies that $\sigma_{1}\approx\sqrt{l}$ and the weight of the regression loss $\frac{1}{2\sigma_{1}^{2}}\approx\frac{1}{2l}$, which is very low. Experimentally, we find that this weight is significantly lower than that of the MLM loss, placing greater emphasis on contextual representation over magnitude-based representation. As a result, we observe poor performance.
To address this issue, we equalized the weights of the two losses and substituted the regression Square Error Loss $\left(y-f_{1}(x,w)\right)^{2}$ with the Logarithmic Square Error Loss (LSEL):
$\left(\log(y+1)-\log\left(f_{1}(x,w)+1\right)\right)^{2}$ (9)
@@
$\mathcal{L}(y)=\frac{1}{2}\textrm{LSEL}+\frac{1}{2}\textrm{MLM}_{loss}$ (10)
This substitution is motivated by the following considerations. Let $\omega$ represent a parameter that requires updating. For simplicity, we’ll denote the partial derivative with respect to $\omega$ as $\partial_{\omega}$. We derive the two following gradients of SEL and LSEL with respect to $\omega$:
$\begin{split}\partial_{\omega}\textrm{SEL}\left(f_{1}(x,w),y\right)&=\partial_{\omega}\left\{\left(y-f_{1}(x,w)\right)^{2}\right\}\\ &=2\partial_{\omega}f_{1}(x,w)\left(f_{1}(x,w)-y\right)\end{split}$ (11)
$\begin{split}\partial_{\omega}\textrm{LSEL}\left(f_{1}(x,w),y\right)&=\partial_{\omega}\left\{\left(\log(y+1)-\log\left(f_{1}(x,w)+1\right)\right)^{2}\right\}\\ &=2\frac{\partial_{\omega}f_{1}(x,w)}{f_{1}(x,w)+1}\log\left(\frac{f_{1}(x,w)+1}{y+1}\right)\end{split}$ (12)
@@
$\frac{\partial_{\omega}\textrm{LSEL}}{\partial_{\omega}\textrm{SEL}}=\frac{1}{\left(f_{1}(x,w)+1\right)\left(f_{1}(x,w)-y\right)}\log\left(\frac{f_{1}(x,w)+1}{y+1}\right)$ (13)
Here, the term $\frac{1}{f_{1}(x,w)-y}\log\left(\frac{f_{1}(x,w)+1}{y+1}\right)$ is contingent on the relative comparison between $f_{1}(x,w)$ and $y$. However, the term $\frac{1}{f_{1}(x,w)+1}$ exhibits a dependency on the magnitude of the prediction. In simpler terms, when high numbers (outliers) are predicted, the Square Error Loss (SEL) will significantly update the weights compared to the Logarithmic Square Error Loss (LSEL), potentially undermining the model’s performance on the real data distribution. This characteristic underscores why LSEL is more effective in handling a wide range of numbers.
Figure  4 provides a concise visual representation of the multi-objective loss mechanism. In this illustration, the first masked token is decoded as the standard word ”Saturation,” assigned the default numerical value of $1$ for non-numerical tokens. The second masked token is decoded as the special token ”NUM,” corresponding to the value $65$.
Figure 4: Brief illustration of the multi objective loss.
VI Experiments

### v3

# Multi-objective Representation for Numbers in Clinical Narratives: A CamemBERT-Bio-Based Alternative to Large-Scale LLMs

@@ Abstract
The processing of numerical values is a rapidly developing area in the field of Language Models (LLMs). Despite numerous advancements achieved by previous research, significant challenges persist, particularly within the healthcare domain. This paper investigates the limitations of Transformers Models in understanding numerical values. Objective: The aim of this research is to categorize numerical values extracted from medical documents into eight specific physiological categories using CamemBERT-bio. Methods: In a context where scalable methods and Large Language Models (LLMs) are emphasized, we explore the option of lifting the limitations of transformer-based models. We examine two strategies: fine-tuning CamemBERT-bio on a small medical dataset with the integration of Label Embedding for Self-Attention (LESA), and combining LESA with additional enhancement techniques such as Xval. Given that CamemBERT-bio is already pre-trained on a large medical dataset, the first approach aims to update its encoder with newly added label embeddings technique, while the second approach seeks to develop multiple representations of numbers (contextual and magnitude-based) to achieve more robust number embeddings. Results: As anticipated, fine-tuning the standard CamemBERT-bio on our small medical dataset did not improve F1 scores. However, significant improvements were observed with CamemBERT-bio + LESA, resulting in an over 13% increase. Similar enhancements were noted when combining LESA with Xval, outperforming conventional methods and giving comparable results to those of GPT-4 Conclusions and Novelty: This study introduces two innovative techniques for handling numerical data, which are also applicable to other modalities. We illustrate how these techniques can improve the performance of Transformer-based models, achieving more reliable classification results even with small datasets.
[…]
Clinical Natural Language Processing, Numerical values classification, Xval, LESA
@@ 1 Introduction
This paper contributes to a broader project aimed at leverage Language Models for diagnosing cardiac failure. According to the National Heart, Lung, and Blood Institute (2023), cardiac failure occurs when the heart is unable to pump sufficient blood to meet the body’s needs, either due to inadequate filling or diminished pumping capacity. Cardiac failure presents through a range of physiological indicators, including blood pressure, ventricular gradients, and ejection fractions. Our research focuses on classifying numerical values extracted from medical notes into one of eight specific physiological categories using CamemBERT-bio [1].
1.1 Goal statement
@@ 1.2 Contribution
  • We assess and juxtapose our tailored model against conventional counterparts including LLMs such as GPT-4. Findings demonstrate that our model surpasses standard approaches and adeptly manages datasets characterized by limited size and imbalance. Additionally, our model delivers performance on par with GPT-4 despite the significant disparity in parameter size.
2 Related works
@@
Processing numerical values with LLMs is a highly active research area. While [4] claimed that most word embedding models could capture numeracy, several limitations persist. Various approaches have been developed to handle numerical values. [5] tackled the issue of number classification by employing a rules-based method utilizing regular expressions. Nevertheless, the success of this approach is largely contingent on the clarity of the text. It depends significantly on consistent abbreviations and the minimal presence of irrelevant textual noise, which can disrupt the regular expression generator’s accuracy. [6] proposed a question-answer pre-training task aimed at improving number understanding. Other studies have explored the potential of different number tokenization methods, such as digit-by-digit, scientific notation, and customized notations [7]. Some previous studies have tackled the challenge with a number-agnostic approach, replacing numbers with placeholder keywords and scaling their embeddings to enforce context-centric encoding [3, 8, 2]. In their work, [2] introduced Label Embedding for Self-Attention (LESA), which incorporates class descriptions directly into the initial feature extraction phase to help the feature extractor focus on relevant features for classification. LESA is similar to EmbedNum [9], which classifies numbers by comparing their embeddings to those of potential categories. However, LESA differs by using similarity exclusively to construct the number’s embedding rather than for the classification process itself. In the current state of the art, advanced LLMs such as LLAMA, GPT-4, Claude are engineered to interpret and adjust to diverse contexts autonomously, making it possible to handle numerical values without requiring any preprocessing.
Recent studies have highlighted several limitations of large language models (LLMs) despite their strong performance across various tasks. While increasing model size has generally led to improved results, some research suggests that LLMs continue to struggle with complex semantic understanding [10]. For instance, [11] found that GPT-3.5 demonstrated high efficiency in extracting critical information from medical notes, but this came at the cost of accuracy, as the model remained prone to errors involving abbreviations and misinterpretations. Similarly, [12] identified significant issues in contextual understanding, interpretability, and bias in training data when evaluating GPT-4 for medical applications. Their findings indicate that these challenges stem from the model’s limited grasp of medical concepts and the insufficient representation of real-world clinical data during training, which could lead to errors and disparities in medical diagnoses. Additionally, the deployment of large-scale LLMs such as GPT-4 raises critical concerns regarding patient data privacy and security, as these models inherently retain vast amounts of training data within their parameters. Given these challenges and the high computational costs associated with LLMs, we prioritize the enhancement of smaller-scale language models, such as CamemBERT-bio, which provide a more adaptable and resource-efficient alternative for medical NLP applications..
@@
Another significant challenge is handling small and imbalanced datasets. While transformer-based architectures generally perform well with large datasets, research by [13, 14, 15] indicates that they may not outperform traditional methods, such as LSTM, when dealing with smaller datasets of clinical notes. A common technique to enhance Language Models performance on small datasets is Knowledge Distillation (KD) [16]. Knowledge Distillation (KD) is based on the principle that a model with fewer parameters is less prone to overfitting. This technique entails training a smaller neural network to replicate the performance of a larger one, using a teacher-student framework. This methodology is the key idea of like DistilBERT [17] and its French counterpart, DistillCamemBERT [18]. Another approach to handling limited data is to create multiple representations for words, leveraging diverse angles of approach to achieve more a robust and generalizable modeling. Typically, this involves constructing an objective loss as the weighted sum of various losses, each specialized in different aspects of the studied modality. However, this approach raises the critical issue of appropriately weighting each loss to avoid bias. In [19], the authors propose an automatic method for weighting multiple losses by optimizing the log-likelihood of the model predictions. Although their approach was applied to images, we found their results relevant and potentially beneficial for our case.
In summary, handling numerical values in NLP remains a challenging and evolving research area. While various approaches, such as rule-based methods, number tokenization techniques, and number-agnostic representations, have been explored, recent advancements in LLMs have enabled more autonomous and context-aware processing of numerical data without requiring extensive preprocessing. However, despite their impressive performance, LLMs still face notable limitations, including challenges in semantic understanding, susceptibility to errors in medical contexts, interpretability issues, and concerns regarding data privacy. Given these challenges, smaller-scale language models, such as CamemBERT-bio, offer a viable alternative by balancing performance, efficiency, and adaptability in specialized domains. Additionally, addressing the challenges of small and imbalanced datasets remains crucial, with techniques like KD and multi-loss optimization proposed to enhance model performance in low-data scenarios.
3 The task and the dataset
@@
4 A few limitations of end-to-end learning for Transformers Models in medical datasets
Most pretrained Transformers Foundational Models such as GPT-4, LLAMA and BERT are pretrained in a end-to-end paradigm. End-to-end learning involves training a model to perform a task from raw input to final output, without intermediate steps or feature engineering. While this is the standard approach for training most Transformers, our work faces the significant limitation of working with small and imbalanced datasets, unlike many state-of-the-art approaches mentioned earlier. This section explores other major limitations of end-to-end training for Transformers Models in the medical domain.
4.1 The Limitations of Tokenization
@@ 4.2 The Structure Issue
As noted in [2], the last excerpt illustrate the lack of contextual information in sentence structures. This excerpt reads a sequence of physiological indicators. The initial reference to FR most likely indicates the respiratory rate (fréquence respiratoire) since a value of 21 is usually too low for the shortening fraction (fraction de raccourcissement). Conversely, the second occurrence of FR refers to a value above 50, suggesting it pertains to the shortening fraction. In this case, it was the values themselves, rather than the context, that provided the necessary distinctions, highlighting a flaw in the sentence structure. This issue echoes the challenges of tokenization, emphasizing the need for a thorough understanding of the numerical values. However, it is worth mentioning that the impact of this issue is very limited on advanced LLMs such as LLAMA, GPT-4, Claude that are engineered to be able to work with data without requiring any preprocessing. Nevertheless, their ability in handling the lack of structure significantly on their prompts [23].
5 Methodology
The methodology adopted in this study involves developing a multi-objective representation of numbers to achieve a more generalizable and robust model [24]. Our representations are obtained through the Masked Language Modelling (MLM) Loss and the Number Prediction Loss. The application of these two representations result in the two following models:
@@ 5 Methodology
5.1 Masked Language Modelling (MLM) objective with LESA
@@
For every individual class, some descriptive keywords are selected, and their initial embeddings are computed and then averaged. This computation produces an embedding matrix, represented as $\boldsymbol{X}^{l}\in\mathbb{R}^{n\times D}$, where $n$ corresponds to the number of classes. This matrix includes unique label embeddings for each class.
@@
In this work, we use this approach to address the limited size of our dataset as it reduces the risks of overfitting while making the model more robust. We then train the model via the Masked Language Modelling (MLM) task as a prefinetuning step. The MLM task involves randomly masking tokens and asking the model to predict the masked tokens, thereby training the model to build contextual representations of tokens.
5.2 Number Prediction Objective
This method consists in four steps as detailed in [3]. Let’s consider an input text $t$ containing both numbers and words. First, all the numbers are replaced by a placeholder keyword NUM, forming a new input text $t_{masked}$. The initial numbers are saved for later. Then, we tokenize and embed $t_{masked}$ resulting in the matrix $h_{masked}$. Subsequently, all the vectors in $h_{masked}$ corresponding to the numbers previously replaced are multiplied by their corresponding value, resulting in a new matrix $h$. Finally, this embeddings matrix $h$ is fed to the encoder of our model. The pipeline is illustrated in Figure  4.
By replacing all numbers by a keyword placeholder, we solve the tokenization limitation mentionned earlier. This loss objective aims at training the model to learn the magnitude of the numbers it encodes. The magnitude information is very important as it is a good discriminator between classification categories. For instance the numbers belonging to heart rate class are usually bigger than those belonging to APGAR class.
5.3 Multi Objective Loss
Working with multiple objectives losses require to merge them into one single loss
@@ 5.3 Multi Objective Loss
where $w_{i}$ represents the weight of the loss $L_{i}$. In our case we have two losses $L_{1}$ which corresponds to the MLM objective and $L_{2}$ which corresponds to the Number Prediction Objective. Our training dataset $\mathcal{D}$ is made of couples $(x,y_{1},y_{2})$ where $x$ is a sequence of tokens with one token masked using the keyword Mask, $y_{1}$ is the index of the masked token in the model vocabulary, and $y_{2}$ is the value of the masked token if it was a number. Otherwise $y_{2}$ is set to $1$. Figure  4 illustrates the inference process of the model. In this illustration, the first masked token is decoded as the standard word ”Saturation,” and the default number $1$ is assigned to it. The second masked token is decoded as the special token ”NUM,” and the number $65$ is predicted by the number prediction head.
Figure 4: Brief illustration of the inference process using multiple representations
In [19], the authors propose an automatic way to compute each loss weight $w_{i}$. For the MLM objective, the predicted token distribution is usually given as a Softmax:
$p(z|f_{1}(x,\Theta),\sigma_{2})=\textrm{Softmax}\left(\frac{1}{\sigma_{1}^{2}}f_{1}(x,\Theta)\right)_{z}$ (3)
@@
where $x$ is the input sequence, $f_{1}(x,\Theta)$ is the vector of logits outputted by the Language Head, $\Theta$ is the model parameters, and $\sigma_{1}$ is artificially added as a temperature parameter. Then, after a few approximations, they get to write the log likelihood as
@@
$\log p(z|f_{1}(x,\Theta))\propto\frac{1}{\sigma_{1}^{2}}\log\textrm{Softmax}\left(f_{1}(x,\Theta)\right)_{z}-\log\sigma_{1}$ (4)
@@
Here, one may recognize the cross entropy loss with its weight $\frac{1}{\sigma_{1}^{2}}$ and a regularization term $\log\sigma_{1}$. Therefore we take $L_{1}$ as the cross entropy loss function and we define $w_{1}=\frac{1}{\sigma_{1}^{2}}$.
For the number prediction task, the authors predict the masked number using a Gaussian distribution:
$p\left(s|f_{2}(x,\Theta)\right)=\mathcal{N}\left(f_{2}(x,\Theta)|\sigma_{2}^{2}\right)$ (5)
where $f_{2}(x,\Theta)$ is Number Prediction Head’s output, and $\mathcal{N}$ refers to the Gaussian probability distribution function and $\sigma_{2}$ its standard deviation. This leads to the log likelihood
$\log p(s|f_{1}(x,\Theta))\propto-\frac{1}{2\sigma_{2}^{2}}\left(s-f_{2}(x,\Theta)\right)^{2}-\log\sigma_{2}$ (6)
Then one can recognize the square error $\left(s-f_{2}(x,\Theta)\right)^{2}$ with its weight $\frac{1}{2\sigma_{2}^{2}}$ and a regularization term $\log\sigma_{2}$. Therefore, we take $L_{2}$ as the square error loss and we define $w_{2}=\frac{1}{2\sigma_{2}^{2}}$.
Finally, when we sum up everything, we get the following loss function:
@@
$\begin{split}L(x,y_{1},y_{2})&=-\frac{1}{\sigma_{1}^{2}}\log\textrm{Softmax}\left(f_{1}(x,\Theta)\right)_{y_{1}}+\frac{1}{2\sigma_{2}^{2}}\left(y_{2}-f_{2}(x,\Theta)\right)^{2}\\ &+\log\sigma_{1}+\log\sigma_{2}\\ &=\frac{1}{\sigma_{1}^{2}}L_{1}(x,y_{1})+\frac{1}{2\sigma_{2}^{2}}L_{2}(x,y_{2})+\log\sigma_{1}+\log\sigma_{2}\end{split}$ (7)
More precisely the global objective function is
$\mathcal{L}(\Theta)=\mathbb{E}_{\mathcal{D}}\left[L(x,y_{1},y_{2})\right]$
with respect to $\Theta,\sigma_{1},\sigma_{2}$ [19]. The first-order conditions with respect to $\sigma_{1}$ and $\sigma_{2}$, give
$\sigma_{1}^{2}=2\mathbb{E}_{\mathcal{D}}\left[L_{1}(x,y_{1})\right]$ (8)
$\sigma_{2}^{2}=\mathbb{E}_{\mathcal{D}}\left[L_{2}(x,y_{2})\right]$ (9)
Since the exact numbers within a medical note are often not predictable from the context, and given the wide range of values present in the dataset (Figure  1), it is natural to assume that $\mathbb{E}_{\mathcal{D}}\left[L_{2}(x,y_{2})\right]$ is significantly larger than $\mathbb{E}_{\mathcal{D}}\left[L_{1}(x,y_{1})\right]$. Therefore by Equations  8 and  9, $w_{2}$ is significantly smaller than $w_{1}$. Our experiments confirms this assumption, placing greater emphasis on contextual representation over magnitude-based representation. As a result, we observe poor performance.
$\mathbb{E}_{\mathcal{D}}\left[L_{2}(x,y_{2})\right]$ depends on the numbers distribution in $\mathcal{D}$ and more precisely on the range of the distribution. In order to lower the impact of this parameter we substituted $L_{2}$ with its log scaled version $\log(L_{2})$. Additionnally, we removed the regularization terms $\log\sigma_{1}$ and $\log\sigma_{2}$ and we adopted $w_{1}=w_{2}=\frac{1}{2}$
$\tilde{L}_{2}(x,y_{2})=\left(\log(y_{2}+1)-\log\left(f_{2}(x,\Theta)+1\right)\right)^{2}$ (10)
@@
$\tilde{L}(x,y_{1},y_{2})=\frac{1}{2}L_{1}(x,y_{1})+\frac{1}{2}\tilde{L}_{2}(x,y_{2})$ (11)
instead of  7.
This substitution is motivated by the following considerations. Let $\theta\in\Theta$ represents a parameter that requires updating. For simplicity, we’ll denote the partial derivative with respect to $\theta$ as $\partial_{\theta}$. We derive the two following gradients of $L_{2}$ and $\tilde{L}_{2}$ with respect to $\theta$:
$\begin{split}\partial_{\theta}L_{2}(x,y_{2})&=\partial_{\theta}\left\{\left(y_{2}-f_{2}(x,\theta)\right)^{2}\right\}\\ &=2\partial_{\theta}f_{2}(x,\theta)\left(f_{2}(x,\theta)-y_{2}\right)\end{split}$ (12)
$\begin{split}\partial_{\theta}\tilde{L}_{2}(x,y_{2})&=\partial_{\theta}\left\{\left(\log(y_{2}+1)-\log\left(f_{2}(x,\theta)+1\right)\right)^{2}\right\}\\ &=2\frac{\partial_{\theta}f_{2}(x,\theta)}{f_{2}(x,\theta)+1}\log\left(\frac{f_{2}(x,\theta)+1}{y_{2}+1}\right)\end{split}$ (13)
@@
$\left\lvert\frac{\partial_{\theta}\tilde{L}_{2}}{\partial_{\theta}L_{2}}\right\rvert=\left\lvert\frac{1}{\left(f_{2}(x,\theta)+1\right)\left(f_{2}(x,\theta)-y_{2}\right)}\log\left(\frac{f_{2}(x,\theta)+1}{y_{2}+1}\right)\right\rvert$ (14)
Here, the term $\frac{1}{f_{2}(x,\theta)-y_{2}}\log\left(\frac{f_{2}(x,\theta)+1}{y_{2}+1}\right)$ is contingent on the relative comparison between the prediction $f_{2}(x,\theta)$ and $y_{2}$. However, the term $\frac{1}{f_{2}(x,\theta)+1}$ exhibits a dependency on the magnitude of the prediction. In simpler terms, when high numbers (outliers) are predicted, $L_{2}$ will significantly update the weights compared to $\tilde{L}_{2}$, potentially undermining the model’s performance on the real data distribution. This characteristic underscores why $\tilde{L}_{2}$ is more effective in handling a wide range of numbers.
6 Experiments